# JAX-MPM: A Learning-Augmented Differentiable Meshfree Framework for GPU-Accelerated Lagrangian Simulation and Geophysical Inverse Modeling

### v1

@@ 5 Inverse modeling of geophysical flows using JAX-MPM
Inverse modeling plays a critical role in geophysical flow analysis, enabling the recovery of unobservable system parameters and guiding the design of effective mitigation strategies. Parameters such as initial velocity fields and spatially varying basal friction fundamentally influence flow dynamics, runout extent, and energy dissipation, yet they are often difficult or impossible to measure directly. In such cases, inverse modeling provides a powerful tool for post-event analysis, model calibration, and data-driven control.
@@
Prior studies have approached inverse problems in geomechanics using techniques ranging from traditional back-analysis [17], Bayesian inference with surrogate models [103], to more recent differentiable simulators [21, 22]. These developments reflect a growing trend toward leveraging gradient-based optimization and automatic differentiation for efficient parameter inference and process learning [35, 84].
@@
In this section, we further evaluate the JAX-MPM framework’s capability for learning-based inverse modeling through a series of parameter estimation tasks that are difficult to address with traditional non-differentiable solvers. Specifically, we consider two representative inverse problems in a dam-break setting: recovering unknown initial velocity fields (Sections 5.15.2) and estimating spatially variable friction coefficients (Section 5.3) from sparse dynamic measurements.
Unless otherwise specified, all inverse modeling experiments in this section are optimized using the Adam algorithm with a decaying learning rate starting at $lr=0.1$.
5.1 Inverse prediction of initial velocity constant
@@
We begin by testing the inverse estimation of a scalar parameter that defines the initial horizontal velocity field in a 2D dam-break scenario (see Section 4.1.1). In this case study, the ground-truth initial velocity is prescribed as $v_{x}^{0}=\overline{\alpha}(H_{0}-y)$ with $\overline{\alpha}=2.0$ and $v_{y}^{0}=0$, where $\overline{\alpha}$ is the unknown parameter to be inferred. The reference simulation assumes an inviscid fluid ($\mu=0$) with a sound speed of $c_{0}=50$ m/s. Both the height and length of the initial water column are $H_{0}=L_{0}=0.5$ m. The computational domain is discretized using a grid spacing of $\Delta h=0.01\,\mathrm{m}$, with four particles per cell. The simulation is run over a duration of $t=0.3$ s with a time step of $\Delta t=3\times 10^{-5}$ s.
@@
The inverse problem is formulated via a PDE-constrained optimization (28), aiming to recover the unknown parameter $\alpha$ by minimizing the loss computed from the discrepancy between simulated dynamics and reference observation data. The objective is evaluated over the final $|\mathcal{T}_{t}|=100$ frames (i.e., the last 0.03 s) sampled from the final 1000 simulation time steps. We consider three observation strategies, each defined by a specific choice of the simulation variable $\bm{z}^{t}$ and spatial indices $\mathcal{T}_{x}$:
@@
(1) Full particle position supervision (Lagrangian): All material points are tracked, with the simulation variable set to particle positions, i.e., $\bm{z}^{t}=\bm{x}^{t}$. Therefore, the observed quantity is defined as $\mathcal{Q}_{l}(\bm{z}^{t})=\bm{x}_{l}^{t}$ for each index $l\in\mathcal{T}_{x}$, where $|\mathcal{T}_{x}|=$ 10,000.
@@
(2) Sparse particle position supervision (Lagrangian): A randomly selected subset of particles is observed, with $|\mathcal{T}_{x}|=100$. The simulation variable remains $\bm{z}^{t}=\bm{x}^{t}$, and the mapping $\mathcal{Q}_{l}(\bm{z}^{t})=\bm{x}_{l}^{t}$ is applied identically.
@@
(3) Sparse velocity-monitor supervision (Eulerian): Predefined spatial regions are used to monitor average particle velocities. The simulation variable is set to particle velocities, $\bm{z}^{t}=\bm{v}^{t}$, and the observed simulated quantity is computed as:
$\mathcal{Q}_{l}(\bm{z}^{t})=\frac{1}{|\mathcal{P}_{l}^{t}|}\sum_{p\in\mathcal{P}_{l}^{t}}\bm{v}_{p}^{t},$ (34)
@@
where $\mathcal{P}_{l}^{t}=\{p\mid\bm{x}_{p}^{t}\in\Omega_{l}\}$ denotes the set of particles located within the $l$-th Eulerian monitor region $\Omega_{l}$. Each $\Omega_{l}$ is defined as a square region with half-size 0.01 m. Here, as shown in Fig. 12, we consider $N_{m}=9$ monitor regions, whose centers are placed on the $3\times 3$ grids located at $\{0.2,0.3,0.4\}\times\{0.1,0.2,0.3\}$.
@@
Optimization starts from an initial guess $\alpha_{0}=0.1$. Fig. 12(a-b) present the training loss and the evolution of the recovered parameter $\alpha^{h}$, respectively. Among the three strategies, velocity supervision achieves the fastest and most stable convergence, despite using the fewest observations. This reflects better alignment between the inferred parameter and the supervised quantity, as $\alpha$ directly influences the velocity field. Interestingly, the sparse particle supervision achieves comparable or even faster convergence than the full-particle case. While the full-field data provides dense information, it also introduces more variability in the gradient signal, which may lead to oscillations during training.
@@
All methods ultimately recover $\alpha$ to values close to the ground truth ($\alpha=2.0$). Fig. 13 visualizes the flow fields at $t=0.15$ s and $t=0.30$ s for the initial guess, the ground truth, and the result obtained using the velocity monitor-based (Eulerian) supervision. The final deposit shape closely matches the reference simulation, indicating that the inferred parameter leads to physically accurate flow dynamics. These results confirm the effectiveness of JAX-MPM in solving inverse problems using both Lagrangian and Eulerian observation types within a unified, fully differentiable framework.
@@
Figure 13: Comparison of flow geometries at (a) $t=0.15$ s and (b) $t=0.30$ s for the initial guess, ground truth, and optimized initial velocity, obtained by using the velocity monitor-based (Eulerian) supervision. The final deposit shape produced by JAX-MPM closely matches the reference solution, confirming the accuracy of scalar parameter recovery. Purple squares indicate the locations of the $N_{m}=9$ velocity monitors, arranged in a $3\times 3$ grid at coordinates $\{0.2,0.3,0.4\}\times\{0.1,0.2,0.3\}$.
5.2 Inverse prediction of initial velocity field
@@
In this example, we adopt the same simulation setup as in the previous experiment (Section 5.1), but consider a more general scenario in which the initial horizontal velocity varies spatially and cannot be described by a single parameter. Specifically, the reference velocity is defined as:
$v_{x}^{0}=2.0\left(1-\left(\frac{y}{H_{0}}\right)^{2}\right)+0.2\sin\left(4\pi\frac{y}{H_{0}}\right),$
@@
which combines a parabolic base profile with a high frequency sine perturbation.
@@
Our objective is to recover the continuous initial velocity field $v_{x}^{0}(y)$ from observational data. To this end, we parameterize the velocity profile using a neural network $\mathcal{N}_{\theta}$, where $\theta$ denotes the trainable weights, and learn a mapping from the vertical coordinate $y$ to the initial velocity $v_{x}^{0}$. Since the goal is to reconstruct a continuous spatial velocity field, velocity monitor data is not employed in this experiment. Instead, we use sparse-particle (Lagrangian) supervision scheme introduced in Section 5.1. To assess the effect of observation sparsity, we randomly select subsets of particles at three sampling densities: 100, 500, and 1000 particles ($|\mathcal{T}_{x}|\in\{100,500,1000\}$). These sampled particle trajectories form the observation dataset. In all cases, the loss is evaluated over the final $|\mathcal{T}_{t}|=100$ frames (last 0.03 s) sampled from the last 1000 time steps, consistent with the previous experiment. We use a feedforward neural network comprising three hidden layers with 30 ReLU-activated neurons each.
@@
As shown in Fig. 14(a–b), the training loss for $|\mathcal{T}_{x}|=500$ and $|\mathcal{T}_{x}|=1000$ decreases steadily and converges after around 400 epochs. In contrast, the loss for $|\mathcal{T}_{x}|=100$ remains nearly flat throughout training, indicating that the data is too sparse to drive meaningful parameter updates. The final relative $L_{2}$ errors between predicted and reference velocity fields are 45.1% for 100 particles, 6.4% for 500, and 3.7% for 1000. These results suggest the presence of an observation sufficiency threshold for effective field reconstruction in this setting: the $|\mathcal{T}_{x}|=100$ case fails to provide adequate information for learning, while increasing the number of observations to 500 and 1000 significantly improves accuracy.
@@
Fig. 14c presents the reconstructed initial velocity profiles against the ground truth. Both the $|\mathcal{T}_{x}|=500$ and $|\mathcal{T}_{x}|=1000$ cases successfully capture the overall profile, particularly in the upper region. The $|\mathcal{T}_{x}|=1000$ result provides a smoother and more accurate reconstruction, while the $|\mathcal{T}_{x}|=500$ case slightly underestimates local variations, due to limited spatial coverage, which constrains the model’s ability to resolve finer features. The deviations near the lower boundary highlight a common challenge in inverse field recovery. This region is influenced by upstream particle motion, making it harder to attribute observed behavior to local initial conditions. Additionally, boundary particles often slow down or stop early due to contact, contributing less to the loss. Together, these factors weaken gradient signals and reduce reconstruction accuracy near the bottom of the domain. In contrast, the $|\mathcal{T}_{x}|=100$ case fails to recover any meaningful spatial structure, resulting in a nearly constant prediction. This highlights again a lower limit on observation density below which continuous field reconstruction becomes ill-posed or ineffective.
These results demonstrate the effectiveness of JAX-MPM for learning spatially varying fields with neural networks. JAX-MPM enables accurate reconstruction of complex initial and boundary conditions via gradient-based optimization, making it a versatile tool for data-driven hazard analysis, model calibration, and predictive simulation.
@@
Figure 14: Inverse recovery of a spatially varying initial velocity field using a neural network. (a) Training loss over epoch, (b) convergence of relative error, and (c) comparison between the learned and reference velocity profiles.
5.3 Inverse prediction of frictional coefficient
@@
Frictional heterogeneity is common in real-world geohazards, and it strongly influences flow mobility, energy dissipation, and deposition patterns. However, frictional properties are rarely measurable in situ, necessitating inverse modeling to recover them from observable outcomes.
[…]
@@
Frictional contact is modeled using the Coulomb friction law proposed by Bardenhagen et al. [6], which modifies the tangential velocity at contact nodes based on particle–wall interaction. The corrected grid velocity $\bm{v}_{i}^{n+1}$ at contact node $i$ is given by:
@@
$\bm{v}_{i}^{n+1}=\tilde{\bm{v}}_{i}^{n+1}-\left[\Delta\bm{v}\cdot\bm{n}_{i}\right]\left(\bm{n}_{i}+\mu^{\prime}\,\bm{n}_{i}\times\bm{\omega}\right),$ (35)
@@
$\Delta\bm{v}:=\tilde{\bm{v}}_{I}^{n+1},\quad\bm{\omega}=\frac{\Delta\bm{v}\times\bm{n}_{I}}{\|\Delta\bm{v}\times\bm{n}_{I}\|},\quad\mu^{\prime}=\min\left[\mu,\frac{\|\Delta\bm{v}\times\bm{n}_{I}\|}{\Delta\bm{v}\cdot\bm{n}_{I}}\right]$ (36)
where $\tilde{\bm{v}}_{i}^{n+1}$ is the intermediate grid velocity before applying contact corrections, and $\bm{n}_{i}$ is the normal vector at the contact surface (e.g., $[0,-1]$ for bottom wall contact).
As shown in Fig. 15, both the $|\mathcal{T}_{x}|=100$ and $|\mathcal{T}_{x}|=500$ cases recover the first three upstream friction segments ($\mu_{0}$, $\mu_{1}$, $\mu_{2}$) with high accuracy, closely matching their ground-truth values. The key difference appears in the downstream segment $\mu_{3}$, where the estimate from $|\mathcal{T}_{x}|=100$ converges to 0.229, corresponding to a 14.5% relative error, while the $|\mathcal{T}_{x}|=500$ case yields a much closer value of 0.198, with only a 1.0% error. This comparison highlights the role of observation density in improving parameter accuracy, particularly for regions that interact with the flow later in time. Additionally, we observe faster convergence on the upstream segments. This reflects the cumulative nature of frictional effects: parameters that influence the flow earlier affect a larger number of particle trajectories and thus produce stronger gradient signals. In contrast, the downstream segments interact with the flow later and more locally, making them harder to recover accurately, especially under sparse supervision.
This example demonstrates that JAX-MPM can accurately recover spatially varying friction coefficients from partial trajectory data, even in contact-dominated systems with discontinuous boundary conditions. Its differentiable Lagrangian formulation enables direct optimization of localized, non-smooth parameters from sparse observations. This capability is essential for real-world geomechanical applications, where frictional heterogeneity is common and observational data is often limited.
Figure 15: Inverse recovery of spatially varying Coulomb friction coefficients in a segmented dam-break domain. (a) Training loss over epoch, (b) Evolution of the predicted friction coefficients $\mu_{f}$ for each segment compared to the ground truth values.

### v2

@@ 5 Inverse modeling of geophysical flows using JAX-MPM
Inverse modeling is essential in geophysical flow analysis, as it enables the recovery of unobservable system parameters and supports the development of effective mitigation strategies. Parameters such as initial velocity fields and spatially varying basal friction [42] fundamentally influence runout extent, flow mobility, and energy dissipation, yet they are often inaccessible through direct measurement. In these contexts, inverse modeling becomes a key tool for post-event analysis, model calibration, and data-informed forecasting.
@@
Prior studies have approached inverse problems in geomechanics using techniques such as traditional back-analysis [16], Bayesian inference with surrogate models [102], and neural network simulators [20, 21]. More recently, there has been growing interest in leveraging gradient-based optimization and algorithmic automatic differentiation to enable efficient parameter inference and process learning in geoscience and solid mechanics applications  [35, 83, 31].
@@
In this section, we further evaluate the JAX-MPM framework’s capability for learning-based inverse modeling through a series of parameter-estimation tasks that are challenging for traditional non-differentiable solvers. Two representative problems in a dam-break setting are considered: (1) recovery of unknown initial velocity fields (Sections 5.15.2) and (2) estimation of spatially variable friction coefficients (Section 5.3) from sparse dynamic measurements. Unless otherwise specified, all inverse modeling experiments in this section are optimized using the Adam algorithm with a decaying learning rate starting at $lr=0.1$.
5.1 Inverse estimation of initial velocity constant
We first test the inverse estimation of a scalar parameter that defines the initial horizontal velocity field in a 2D dam-break scenario (see Section 4.1.1). The ground-truth initial velocity field is prescribed as $v_{x}^{0}=\overline{\alpha}(H_{0}-y)$ and $v_{y}^{0}=0$, with $\overline{\alpha}=2.0$. For the inverse modeling task, $\overline{\alpha}$ is treated as an unknown coefficient to be inferred.
@@ 5.1 Inverse estimation of initial velocity constant
The reference simulation model assumes an inviscid fluid ($\mu=0$) with a numerical sound speed of $c_{0}=50$ m/s. The initial water column has height and length $H_{0}=L_{0}=0.5$ m, discretized using a grid spacing of $\Delta h=0.01\,\mathrm{m}$ with four particles per cell. The simulation is run for a total duration of $t=0.3$ s with a fixed time step $\Delta t=3\times 10^{-5}$ s.
@@
The inverse problem is formulated as a PDE-constrained optimization (30), where the objective is to minimize the discrepancy between simulated dynamics and reference observations over the final $|\mathcal{T}_{t}|=100$ frames, sampled every 10 steps from the last 1000 simulation steps (i.e., the final 0.03 s of the simulation).
We consider three observation strategies, each characterized by a specific choice of simulation variable $\bm{z}^{t}$ and spatial observation indices $\mathcal{T}_{x}$:
@@
Case 1 - Full particle position supervision (Lagrangian): All material points are tracked, with the observation variable defined as particle positions, i.e., $\bm{z}^{t}=\bm{x}^{t}$. The observed quantity is then $\mathcal{Q}_{l}(\bm{z}^{t})=\bm{x}_{l}^{t}$ for each index $l\in\mathcal{T}_{x}$, with $|\mathcal{T}_{x}|=$ 10,000.
@@
Case 2 - Sparse particle position supervision (Lagrangian): A randomly selected subset of $|\mathcal{T}_{x}|=100$ particles is tracked, with the same observation mapping as in the full supervision Case 1.
@@
Case 3 - Sparse velocity-monitor supervision (Eulerian): Predefined spatial regions are used to monitor average particle velocities. The simulation variable is set to the particle velocities, $\bm{z}^{t}=\bm{v}^{t}$, and the observed simulated quantity is computed as:
$\mathcal{Q}_{l}(\bm{z}^{t})=\frac{1}{|\mathcal{P}_{l}^{t}|}\sum_{p\in\mathcal{P}_{l}^{t}}\bm{v}_{p}^{t},\quad\mathcal{P}_{l}^{t}=\{p\mid\bm{x}_{p}^{t}\in\Omega_{l}\}$ (37)
@@
where $\mathcal{P}_{l}^{t}$ denotes the set of particles located within the $l$-th Eulerian monitor region $\Omega_{l}$. For example, as illustrated in Fig. 13, we consider $N_{m}=9$ monitor regions, each region is a square with a half-side length of 0.01 m. The centers of these regions are arranged in a $3\times 3$ grid at spatial locations $\{0.2,0.3,0.4\}\text{ m}\times\{0.1,0.2,0.3\}\text{ m}$.
@@
The JAX-MPM optimization begins from an initial guess $\alpha_{0}=0.1$. Fig. 12 presents the evolution of both the training loss and the estimated parameter $\alpha^{h}$ over the course of learning iterations, demonstrating that all three observation strategies successfully converge to the ground truth ($\overline{\alpha}=2.0$). Among them, velocity-monitor supervision achieves the fastest and most stable convergence, despite using the fewest observations. This reflects a strong alignment between the supervised quantity and the underlying parameter, as $\alpha$ directly governs the velocity field. Notably, the sparse particle supervision (Case 2) achieves convergence comparable to—or even faster than—that of full particle supervision (Case 1). We conjecture that, while full-field observations provide denser information, they may also introduce greater complexity into the loss landscape and its gradients, resulting in increased oscillations during optimization, particularly when the underlying parameter is spatially uniform.
@@
Fig. 13 visualizes the flow fields at $t=0.15$ s and $t=0.30$ s for the initial guess, the ground truth, and the optimized result obtained using the velocity monitor-based (Eulerian) supervision. The final deposit shape closely matches the reference simulation, indicating that the inferred parameter leads to physically accurate flow dynamics. These results confirm the effectiveness of JAX-MPM in solving inverse problems using both Lagrangian and Eulerian observation types within a unified, fully end-to-end differentiable framework.
@@
Figure 13: Comparison of flow geometries at (a) $t=0.15$ s and (b) $t=0.30$ s for the initial guess, ground truth, and optimized initial velocity, obtained by using the velocity monitor-based (Eulerian) supervision. Purple squares indicate the locations of the $N_{m}=9$ velocity monitors.
5.2 Inverse estimation of initial velocity field
@@
We extend the previous example to a more general scenario in which the initial horizontal velocity field varies spatially and cannot be represented by a single scalar parameter. The reference horizontal velocity field is defined as
$v_{x}^{0}(y)=2.0\left(1-\left(\frac{y}{H_{0}}\right)^{2}\right)+0.2\sin\left(4\pi\frac{y}{H_{0}}\right),$ (38)
@@
which combines a parabolic base profile with a sinusoidal perturbation.
The objective is to recover the continuous velocity field $v_{x}^{0}(y)$ from the observational data. To this end, we parametrize the field using a feedforward neural network $\mathcal{N}_{\theta}$ (c.f. Eq. 34), consisting of three hidden layers with 30 ReLU-activated units each. The network maps the vertical coordinate $y$ to the corresponding velocity $v_{x}^{0}(y)$, where $\theta$ denotes the trainable weights.
@@
Since the goal is to reconstruct a continuous spatial velocity field, velocity monitor data are excluded in this experiment to better showcase the applicability of the proposed framework. Instead, we adopt the sparse particle (Lagrangian) position supervision scheme introduced in Section 5.1. To assess the effect of observation sparsity, we randomly sample particles at three resolutions: 100, 500, and 1000 particles ($|\mathcal{T}_{x}|\in\{100,500,1000\}$). The trajectories of these sampled particles constitute the observation dataset. In all cases, the loss is evaluated over the final $|\mathcal{T}_{t}|=100$ frames (last 0.03 s) sampled from the last 1000 time steps, which is consistent with the experiment setup in Section 5.1.
@@
As shown in Fig. 14(a–b), the training loss for $|\mathcal{T}_{x}|=500$ and $|\mathcal{T}_{x}|=1000$ decreases steadily and converges after around 400 epochs. In contrast, the loss for $|\mathcal{T}_{x}|=100$ remains nearly flat throughout training, indicating that the observation density is insufficient to drive meaningful parameter updates. The final relative $L_{2}$ errors between predicted and reference velocity fields are 45.1% for 100 observation particles, 6.4% for 500, and 3.7% for 1000. These results suggest there could exist a threshold in observation density for effective field reconstruction in this setting.
@@
Fig. 14c compares the reconstructed initial velocity profiles against the ground truth. Both the $|\mathcal{T}_{x}|=500$ and $|\mathcal{T}_{x}|=1000$ cases accurately capture the overall profile, especially in the upper region. The $|\mathcal{T}_{x}|=1000$ case provides a smoother and more accurate reconstruction, while the $|\mathcal{T}_{x}|=500$ case slightly underestimates local variations due to limited spatial coverage, which restricts the model’s ability to resolve finer features. In contrast, the $|\mathcal{T}_{x}|=100$ case fails to recover any meaningful spatial structure, resulting in a nearly constant initial velocity prediction. This again underscores the importance of sufficient observation density for successful continuous field reconstruction.
Deviations near the lower boundary (i.e., close to $y=0$) reveal a common challenge in inverse field recovery (see Fig. 14c). This region is influenced by complex particle dynamics arising from both upstream motion and contact interactions, making it more difficult to attribute observed behavior to local velocity features. As a result, parameter reconstruction becomes relatively ill-posed and tends to degrade near the base of the domain. Nonetheless, the global approximation capability of the neural network parameterization maintains robust estimation even in this challenging region.
These results demonstrate the effectiveness of JAX-MPM in learning spatially varying fields with neural-network-based representations.
@@
Figure 14: Inverse recovery of a spatially varying initial velocity field using a neural network. (a) Training loss over epochs, (b) convergence of relative $L_{2}$ error, and (c) comparison between the learned and reference velocity profiles.
5.3 Inverse estimation of frictional coefficient
@@
Frictional heterogeneity is prevalent in real-world geohazard scenarios and significantly affects flow mobility and deposition patterns. However, frictional properties are rarely measurable in situ, making inverse modeling an essential tool for estimating such parameters from observable outcomes.
In this example, we demonstrate the ability of JAX-MPM to recover spatially varying Coulomb friction coefficients $\mu_{f}$ along the bottom boundary of a 2D dam-break scenario using particle-trajectory data.
The simulation is initialized in a standard dam-break configuration within a domain of length $L=1.6\,\mathrm{m}$ and height $H=0.4\,\mathrm{m}$, with an initial water column of $L_{0}=0.5\,\mathrm{m}$ and $H_{0}=0.3\,\mathrm{m}$. The bottom surface is divided into four equal-length segments, each assigned a ground-truth friction coefficient, i.e., $(\mu_{0},\mu_{1},\mu_{2},\mu_{3})=(0.0,0.5,0.1,0.2)$, which are treated as trainable parameters during optimization. Since each segment interacts with the flow at different times, supervision is provided using particle trajectories sampled from $|\mathcal{T}_{t}|=1000$ observation frames, uniformly distributed over the full 10,000-step simulation (corresponding to 0.4 s with a time step of $\Delta t=4\times 10^{-5}\,\mathrm{s}$). Two levels of observation sparsity are tested: $|\mathcal{T}_{x}|=100$ and $|\mathcal{T}_{x}|=500$ randomly selected particles.
@@
Frictional contact is modeled using the Coulomb friction law, which modifies the tangential velocity at contact nodes based on particle–wall interaction. The corrected grid velocity $\tilde{\bm{v}}_{i}^{n+1}$ at contact node $i$ is given by [6]:
@@
$\tilde{\bm{v}}_{i}^{n+1}=\bm{v}_{i}^{n+1}-\left[\Delta\bm{v}\cdot\bm{n}_{i}\right]\left(\bm{n}_{i}+\mu^{\prime}\,\bm{n}_{i}\times\bm{\omega}\right),$ (39)
@@
$\Delta\bm{v}:=\bm{v}_{i}^{n+1}-\bm{v}^{\text{cm}}_{i},\quad\bm{\omega}=\frac{\Delta\bm{v}\times\bm{n}_{i}}{\|\Delta\bm{v}\times\bm{n}_{i}\|},\quad\mu^{\prime}=\min\left[\mu,\frac{\|\Delta\bm{v}\times\bm{n}_{i}\|}{\Delta\bm{v}\cdot\bm{n}_{i}}\right]$ (40)
where $\bm{v}^{\text{cm}}_{i}$ is the center-of-mass velocity of the two bodies at contact node $i$, which reduces to zero as one of the bodies is a stationary rigid wall [34]. $\mu$ is the friction coefficient, $\bm{v}_{i}^{n+1}$ is the intermediate grid velocity before contact correction, $\bm{n}_{i}$ is the normal vector at the contact surface (e.g., [0,1][0,-1] for bottom wall contact), $\bm{\omega}$ is the unit tangential vector, and $\mu^{\prime}$ is the effective friction coefficient.
Fig. 15b shows the evolution of the four estimated friction coefficients under the two different observation densities throughout the optimization process. The first three upstream segments ($\mu_{0},\mu_{1},\mu_{2}$) are accurately identified in both cases. The most notable discrepancy appears in the downstream segment $\mu_{3}$: with $|\mathcal{T}_{x}|=100$, the estimated coefficient converges to 0.229 (14.5% error), while with $|\mathcal{T}_{x}|=500$, it converges to 0.198 (1.0% error). This highlights the role of observation density in improving parameter accuracy, particularly for downstream regions that interact less significantly with the flow.
The convergence behavior reflects the cumulative nature of frictional effects. Upstream parameters, which influence a larger number of particle trajectories, generate stronger gradient signals and thus converge more rapidly. In contrast, downstream parameters interact with the flow later and more locally, resulting in weaker gradients, especially under sparse supervision, making them more challenging to accurately recover.
Figure 15: Inverse recovery of spatially varying Coulomb friction coefficients in a segmented dam-break domain. (a) Training loss over epochs, (b) Evolution of the predicted friction coefficients $\mu$ for each segment compared to the ground-truth values.
This example demonstrates that JAX-MPM can reliably recover spatially varying friction coefficients from partial trajectory data, even in contact-dominated systems with discontinuous boundary conditions. Its differentiable Lagrangian formulation enables direct optimization of localized, non-smooth parameters from sparse observations. This capability is essential for real-world geomechanical applications, where frictional heterogeneity is common and observational data are often limited.